Shift Summary 27 June, 1999, Owl

Jim Hirschauer (jhirsch@cuhep.Colorado.EDU)
Sun, 27 Jun 1999 06:06:09 -0600 (MDT)

Shift Summary for owl shift, 27/Jun/99

On Shift: Theo Alexopoulos, Jim Hirschauer

Hours of beam delivered at all intensities: 6.5
Hours of beam delivered at nominal intensity: 6.5
Hours of E832 data (special runs not included): 6.75
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: .25 hrs - starting new run

========================================================================
Runs ended this shift: 13744
Run 13744 - 11.6M events - BEAM_832
Run 13745 - in progress - BEAM_832

Tapes brought to FCC this shift: QKC046,QKC047,QKC048
QKS046,QKS047,QKS048
QKK046,QKK047,QKK048
QKB017

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
Smooth E832 datataking.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13744 00:15 mounted tapes QKPE03 and QKNL03 in drives
ktev3,tps2d3 and ktev2,tps3d2, respectively,
as requested by onsplit

60 00:30 JAWS noticed vertical beam position for SEED2
was completely off (at -4) and called MCR.
They were trying to change something in the
TeV to improve the tune. They went back to the
original configuration

64 beam back on target

01:45 mounted tape QKN004 into drive ktev1,tps3d2, as
requested by onsplit

02:30 in histogram trg#summary, "Level 3 CPU Summary:" one
processor seemed to be above nominal limits, no
correction made

03:30 had ktevana errors, ran kfrend

04:45 saw problems with HA, channel 28 is below nominal
limits for ADC mean value

236 04:46 Scaler sdaq gave warning that CA_left/right were
were well below nominal limits. CA high voltage
was fine and SEM was reading 3E12. Online histograms
showed no counts in the CA's. TheoA noticed the drift
chambers were drawing no current.
The beam stop was found to be in and the MCR was
called. The MCR determined that NM2V was not ramping
properly and tripped the beam stop.

251 05:00 With NM2V ramping correctly, the beam stop was removed
and normal data taking resumed. Spills 236-251 were
affected.

05:10 received ktevana errors, ran kfrend

05:33 Ended Run 13744
Mounted new tapes

13745 05:45 Started Run 13745 in BEAM_832 mode, NM4AN<0

06:30 Linac Vacuum burst; Est to beam: 10mins
07:00 Still no beam, new est 20-25 mins

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

o keep collecting BEAM_832 data